# Optimal Spectral Transitions in High-Dimensional Multi-Index Models

### v1

@@ A.2 State Evolution of Linear GAMP
that characterize respectively the alignment between $\bm{\mathbf{\hat{W}}}^{t}$ and the weights $\bm{W}_{\star}$ and the norm of $\bm{\mathbf{\hat{W}}}^{t}$. In this appendix we present the state evolution equations for the GAMP algorithm (60, 61) with linear denoiser functions, while we refer to Aubin et al. [2018] for a complete derivation in more general settings:
@@
$\hat{{\bm{M}}}^{t}$ $=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}({\mathrm{y}}){\bm{M}}^{t}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}})\right],$ (67)
$\hat{{\bm{Q}}}^{t}$ $=\alpha\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}({\mathrm{y}}){\bm{M}}^{t}(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}}){\bm{M}}^{T}\bm{G}({\mathrm{y}})^{T}\right]+\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}({\mathrm{y}}){\bm{Q}}^{t}\bm{G}({\mathrm{y}})^{T}\right]\right).$ (68)
Appendix B Derivation of the main results - Asymmetric spectral method
@@
$\bm{\Omega}^{t}$ $=\gamma^{-1}{\underline{\bm{L}}}\cdot\bm{\Omega}^{t-1},$ (69)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\gamma^{-1}\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t},$ (70)
with ${\underline{\bm{G}}}$ defined in eq. (14). Given the principal eigenpair $\gamma_{1},\bm{\Omega}_{1}$ of ${\underline{\bm{L}}}$, the spectral estimator defined in Definition 16 is a fixed point of this algorithm for $\gamma=\gamma_{1}$ and $\bm{\Omega}\propto\bm{\Omega}_{1}\frac{\gamma_{1}\sqrt{d}}{||\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}_{1}||}$. Interestingly, the above algorithm is a linear GAMP, with denoiser functions ${\bm{g}}_{\rm out}^{t}(\bm{\omega},y)=(\mathrm{Cov}[\bm{\mathbf{z}}|y]-{\bm{I}})\bm{\omega}$ and $\bm{f}_{in}^{t}(\bm{b})=\gamma^{-1}\bm{b}$, $\forall t$, and Onsager terms
${\bm{A}}_{t}\xrightarrow{n\to\infty}$ $\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}\big{|}{\mathrm{y}}]-\alpha{\bm{I}}={\bm{0}},$ (71)
as already shown in (64), and ${\bm{V}}_{t}=\gamma^{-1}{\bm{I}}$. In fact, we can rewrite the generalized power iteration algorithm as
$\bm{\Omega}^{t}$ $=\gamma^{-1}(\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}-{\bm{I}}_{n}){\underline{\bm{G}}}\cdot\bm{\Omega}^{t-1}=\bm{\mathbf{X}}(\gamma^{-1}\bm{B}^{t})-\gamma^{-1}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t-1},$ (72)
$\bm{B}^{t+1}$ $=\bm{\mathbf{X}}^{T}{\underline{\bm{G}}}\cdot\bm{\Omega}^{t},$ (73)
@@
This algorithm does not offer any particular advantage for the practical computation of the spectral estimator compared to other algorithms for estimating the principal eigenvector. However, as an Approximate Message Passing algorithm, it enables to track low-dimensional functions of the iterate $\bm{\mathbf{\hat{W}}}^{t}$ via the associated state evolution. Specifically, we will analyze the weak recovery properties of the spectral method by studying the convergence of the state evolution equations
@@
where
${\mathcal{F}}(\bm{M})$ $\coloneqq\mathbb{E}_{y}[(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})\bm{M}(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})],$ (77)
${\mathcal{G}}(\bm{M})$ $\coloneqq\mathbb{E}_{y}[(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})\bm{M}(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})\bm{M}^{T}(\mathrm{Cov}[\bm{\mathbf{z}}\big{|}y]-{\bm{I}})].$ (78)
@@
    Informed fixed points. For $\gamma=\alpha\nu_{k}$ ($\nu_{k}\neq 0$), $\bm{M}\propto\bm{M}_{k}$, is a non-zero fixed point for eq. (75). In particular, we are interested to the fixed point correspondent to the largest $\gamma$, i.e. ${\bm{M}}\propto{\bm{M}}_{1}$. For $\alpha>\alpha_{c}$ and $\gamma=\nicefrac{{\alpha}}{{\alpha_{c}}}>1$, the largest eigenvalue of $\alpha{\mathcal{F}}(\cdot)$ is equal to one and any ${\bm{M}}\propto{\bm{M}}_{1}$ is a stable fixed point. Moreover, eq. (76) at convergence:
    $\bm{Q}=\bm{M}^{2}+\frac{\alpha_{c}^{2}}{\alpha}(m^{2}{\mathcal{G}}(\bm{M})+{\mathcal{F}}(\bm{Q}))\implies$ (79)
@@
    Therefore, the correspondent estimator weakly recovers $\bm{W}_{\star}$.
@@
    Uninformed fixed point. Initializing GAMP in a subspace orthogonal to the signal, i.e. $\bm{M}^{0}={\bm{0}}$,
    $\bm{Q}=\gamma^{-2}\alpha{\mathcal{F}}(\bm{Q}),$ (83)
    therefore, for $\gamma=\sqrt{\alpha\nu_{1}}$, $\bm{M}={\bm{0}}$, $\bm{Q}=\bm{M}_{1}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$ is a fixed point of the state evolution. Note that, since the largest eigenvalue of $\alpha{\mathcal{F}}(\cdot)$ is larger than one, such fixed point is unstable. Since the proposed GAMP is equivalent to a generalized power iteration algorithm, normalized by the constant $\gamma\in\mathbb{R}$, it converges only if $\gamma$ corresponds to the absolute value of the eigenvalue with largest magnitude in the subspace of initialization.
Appendix C Derivation of the main results - Symmetric spectral method
Similarly to what we have done in the previous section, we introduce a GAMP algorithm that will serve as a framework to study the properties of the spectral estimator defined in (17).
Definition C.1.
@@ Appendix C Derivation of the main results - Symmetric spectral method
${\bm{g}}_{\rm out}^{t}({\mathrm{y}},\bm{\omega})=\bm{\mathcal{T}}(y){\bm{V}}_{t}^{T}(a_{t}{\bm{I}}-{\bm{V}}_{t}\bm{\mathcal{T}}(y){\bm{V}}_{t}^{T})^{-1}\bm{\omega},\quad\quad\bm{f}_{in}^{t}({\bm{b}})=(\gamma_{t}{\bm{V}}_{t}^{T}-{\bm{A}}_{t})^{-1}{\bm{b}},$ (84)
@@
Dropping the time index for the fixed-point variables and parameters, and defining ${\underline{\bm{T}}}_{{\bm{V}}}$ as
${\underline{\bm{T}}}_{{\bm{V}}}\cdot\bm{\Omega}$ $={\bm{V}}{\underline{\bm{T}}}\cdot\bm{\Omega},$ (85)
the fixed points satisfy
$\bm{\Omega}=\bm{X}\bm{\mathbf{\hat{W}}}-{\underline{\bm{T}}}_{{\bm{V}}}\cdot\bm{\Omega}\implies\bm{\Omega}=\left({\underline{\bm{I}}}+{\underline{\bm{T}}}_{{\bm{V}}}\right)^{-1}\cdot(\bm{X}\bm{\mathbf{\hat{W}}}),$ (86)
$\bm{\mathbf{\hat{W}}}(\gamma_{t}{\bm{V}}-{\bm{A}}^{T})=\bm{\mathbf{X}}X^{T}{\underline{\bm{T}}}\cdot\left(\left({\underline{\bm{I}}}+{\underline{\bm{T}}}_{{\bm{V}}}\right)^{-1}\cdot(\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}})\right)-\bm{\mathbf{\hat{W}}}{\bm{A}}^{T}\implies a\gamma_{t}\bm{\mathbf{\hat{W}}}{\bm{V}}={\underline{\bm{T}}}\cdot(\bm{\mathbf{\hat{W}}}{\bm{V}}),$ (87)
@@
$[{\underline{\bm{T}}}\cdot(\left({\underline{\bm{I}}}+{\underline{\bm{T}}}_{{\bm{V}}}\right)^{-1}\cdot\bm{\Omega})]_{i.}$ $=\bm{\mathcal{T}}(y){\bm{V}}^{T}(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T})^{-1}[\left({\underline{\bm{I}}}+{\underline{\bm{T}}}_{{\bm{V}}}\right)^{-1}\bm{\Omega}]_{i.}$ (88)
$=\bm{\mathcal{T}}(y){\bm{V}}^{T}(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T})^{-1}\left({\bm{I}}+{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T}(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T})^{-1}\right)^{-1}\bm{\Omega}_{i.}$ (89)
$=\bm{\mathcal{T}}(y){\bm{V}}^{T}\left(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T}+{\bm{V}}\bm{\mathcal{T}}(y){\bm{V}}^{T}\right)^{-1}\bm{\Omega}_{i.}$ (90)
$=a^{-1}\bm{\mathcal{T}}(y){\bm{V}}^{T}\bm{\Omega}_{i.}$ (91)
Therefore, $\bm{\mathbf{\hat{W}}}=\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}{\bm{V}}^{-1}$ is a fixed point of the algorithm, for $a_{t}$ and $\gamma_{t}$ appropriately chosen, with eigenvalue given by $a_{t}\gamma_{t}$ at convergence. [^2] Note that the overlap matrices for this algorithm refer to $\bm{\mathbf{\hat{W}}}_{{\underline{\bm{T}}}}{\bm{V}}^{-1}$ and not directly to the spectral estimator itself. However, if $\|{\bm{M}}\|_{\mathrm{F}}>0$, the weak recovery condition is satisfied.

### v2

@@ A.2 State Evolution of Linear GAMP
that characterize respectively the alignment between $\bm{\mathbf{\hat{W}}}^{t}$ and the weights $\bm{W}_{\star}$ and the norm of $\bm{\mathbf{\hat{W}}}^{t}$. In this appendix we present the state evolution equations for the GAMP algorithm (42, 44) with linear denoiser functions, while we refer to Aubin et al. [2018] for a complete derivation in more general settings:
@@
$\hat{{\bm{M}}}^{t}$ $=\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}_{\rm out}({\mathrm{y}}){\bm{M}}^{t}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]\right],$ (50)
$\hat{{\bm{Q}}}^{t}$ $=\alpha\left(\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}_{\rm out}({\mathrm{y}}){\bm{M}}^{t}\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]{\bm{M}}^{T}\bm{G}_{\rm out}({\mathrm{y}})^{T}\right]+\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\bm{G}_{\rm out}({\mathrm{y}}){\bm{Q}}^{t}\bm{G}_{\rm out}({\mathrm{y}})^{T}\right]\right).$ (51)
Appendix B Proof Outlines for Theorems 3.3 and 3.4 – Asymmetric Spectral Method
@@
$\bm{\omega}^{t}$ $=\gamma^{-1}\bm{L}\,\bm{\omega}^{t-1}\in\mathbb{R}^{np},$ (52)
$\bm{\mathbf{\hat{W}}}^{t+1}$ $=\gamma^{-1}\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{\bm{G}}\,\bm{\omega}^{t}\right)\in\mathbb{R}^{d\times p},$ (53)
with $\hat{\bm{G}}$ defined in eq. (9) and $\gamma$ a parameter to be fixed. A pair $\bm{\omega}\neq{\bm{0}}$, $\bm{\mathbf{\hat{W}}}=\gamma^{-1}\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{{\bm{G}}}\bm{\omega}\right)$, is a fixed point of the above algorithm if and only if ${\bm{L}}\bm{\omega}=\gamma\bm{\Omega}$. Thus, given the largest real $\gamma$ such that the algorithm has a non-zero fixed point, the latter will correspond to the asymmetric spectral estimator in Definition 10. Interestingly, the above algorithm is also linear GAMP A.1, with denoiser functions ${\bm{g}}_{\rm out}^{t}(\bm{\omega}\in\mathbb{R}^{p},y)=\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}=y]\bm{\omega}$ and $\bm{f}_{in}^{t}(\bm{b})=\gamma^{-1}\bm{b}$, $\forall t$, and Onsager terms
${\bm{A}}_{t}\xrightarrow{n\to\infty}$ $\alpha\mathbb{E}_{{\mathrm{y}}\sim\mathsf{Z}}\left[\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}\big{|}{\mathrm{y}}]\right]={\bm{0}}$ (54)
(as shown in (47)), and ${\bm{V}}_{t}=\gamma^{-1}{\bm{I}}$. In fact, we can rewrite the generalized power iteration algorithm as in Definition 3.1
$\bm{\Omega}^{t}$ $=\gamma^{-1}(\bm{\mathbf{X}}\bm{\mathbf{X}}^{T}-{\bm{I}}_{n}){\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\bm{\Omega}^{t-1}\right)\right)=\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}^{t}-\gamma^{-1}{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\bm{\Omega}^{t-1}\right)\right),$ (55)
$\bm{B}^{t+1}$ $=\bm{\mathbf{X}}^{T}{\rm mat}\left(\hat{\bm{G}}\,{\rm vec}\left(\bm{\Omega}^{t}\right)\right),$ (56)
@@
B.1 State Evolution
As an Approximate Message Passing algorithm, this algorithm enables to track low-dimensional functions of the iterate $\bm{\mathbf{\hat{W}}}^{t}$ via the associated state evolution. Specifically, we will analyze the weak recovery properties of the asymmetric spectral method by studying the convergence of the state evolution equations A.2:
@@ B.1 State Evolution
where, recalling the notation ${\bm{G}}({\mathrm{y}})=\mathbb{E}[\bm{\mathbf{z}}\bm{\mathbf{z}}^{T}-{\bm{I}}|{\mathrm{y}}]$,
${\mathcal{F}}(\bm{M})$ $\coloneqq\mathbb{E}_{y}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})],$ (60)
${\mathcal{G}}(\bm{M})$ $\coloneqq\mathbb{E}_{y}[{\bm{G}}({\mathrm{y}})\bm{M}{\bm{G}}({\mathrm{y}})\bm{M}^{T}{\bm{G}}({\mathrm{y}})].$ (61)
@@
    Informed fixed points. (Theorem 3.3) $\forall\alpha$, for $\gamma=\alpha\nu_{k}$ ($\nu_{k}\neq 0$), $\bm{M}\propto\bm{M}_{k}$, is a non-zero fixed point for eq. (58). In particular, since the asymmetric spectral estimator is the fixed point correspondent to the largest $\gamma$, we are interested to the case ${\bm{M}}\propto{\bm{M}}_{1}$. We show now that $\|{\bm{M}}\|_{\mathrm{F}}\neq 0$ (i.e. the fixed point is actually informative) for $\alpha>\alpha_{c}$. Given $\gamma=\nicefrac{{\alpha}}{{\alpha_{c}}}>1$, the largest eigenvalue of the operator $\gamma^{-1}\alpha{\mathcal{F}}(\cdot)$ is equal to 1 and any ${\bm{M}}\propto{\bm{M}}_{1}$ is a stable fixed point. Moreover, eq. (59) at convergence:
    $\bm{Q}=\bm{M}^{2}+\frac{\alpha_{c}^{2}}{\alpha}(\|{\bm{M}}\|_{\mathrm{F}}^{2}{\mathcal{G}}(\bm{M})+{\mathcal{F}}(\bm{Q}))\implies$ (62)
@@
    Therefore, the correspondent estimator $\bm{\mathbf{\hat{W}}}$ eq. (57) weakly recovers $\bm{W}_{\star}$.
    Note that, although it is beyond the scope of this work, the presented framework enables the identification of additional informative eigenvectors (with real eigenvalues smaller than $\alpha/\alpha_{c}$), emerging from the bulk at larger sample complexities.
@@
    Uninformed fixed point. (Theorem 3.4) Initializing GAMP in a subspace orthogonal to the signal, i.e. $\bm{M}^{0}={\bm{0}}$, we have that ${\bm{M}}^{t}={\bm{0}}$ at all times and eq. (59) at convergence
    $\bm{Q}=\gamma^{-2}\alpha{\mathcal{F}}(\bm{Q}).$ (66)
    Therefore, for $\gamma=\sqrt{\alpha\nu_{1}}$, $\bm{M}={\bm{0}}$, $\bm{Q}=\bm{M}_{1}\in\mathbb{S}_{+}^{p}\setminus\{{\bm{0}}\}$ is a fixed point of the state evolution. Note that, since the largest eigenvalue of $\gamma^{-1}\alpha{\mathcal{F}}(\cdot)$ in eq. (58) is $\sqrt{\alpha/\alpha_{c}}$, ${\bm{M}}={\bm{0}}$ is a stable fixed point for $\alpha\leq\alpha_{c}$, and unstable otherwise. Since the proposed GAMP is a generalized power iteration algorithm, normalized by the constant $\gamma\in\mathbb{R}$, it converges only if $\gamma$ corresponds to the absolute value of the eigenvalue with largest magnitude in the subspace of initialization.
Appendix C Proof Outlines for Theorems 3.8 and 3.9 – Symmetric Spectral Method
Similarly to what we have done in the previous section, we introduce a GAMP algorithm that will serve as a framework to study the properties of the spectral estimator defined in (12). We stress that this algorithm does not offer any particular advantage for the practical computation of the spectral estimator compared to other spectral algorithms.
@@
${\bm{g}}_{\rm out}^{t}({y},\bm{\omega})=\bm{\mathcal{T}}(y){\bm{V}}_{t}^{T}(a_{t}{\bm{I}}-{\bm{V}}_{t}\bm{\mathcal{T}}(y){\bm{V}}_{t}^{T})^{-1}\bm{\omega},\quad\quad\bm{f}_{in}^{t}({\bm{b}})=(\gamma_{t}{\bm{V}}_{t}^{T}-{\bm{A}}_{t})^{-1}{\bm{b}},$ (67)
@@
We first show that, for suitable choiches of $a_{t}$ and $\gamma_{t}$, the non-zero fixed point of this algorithm correspond to an eigenvector of ${\bm{T}}$. Dropping the time index for the fixed-point variables and parameters, and defining $\hat{{\bm{G}}}_{{\bm{T}}}\in\mathbb{R}^{np\times np}$ as
$\left[\hat{{\bm{G}}}_{\bm{\mathcal{T}}}\right]_{(i\mu),(j\nu)}\coloneqq\delta_{ij}[\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}\left(a_{t}{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}\right)^{-1}]_{\mu\nu},\quad i,j\in\llbracket n\rrbracket,\,\mu\nu\in\llbracket p\rrbracket,$ (68)
for $i\in\llbracket n\rrbracket,\,\mu\in\llbracket p\rrbracket$, the fixed points satisfy
$\Omega_{i\mu}=[\bm{X}\bm{\mathbf{\hat{W}}}]_{i\mu}-\sum_{j\in\llbracket n\rrbracket}\sum_{\nu\in\llbracket p\rrbracket}\underbrace{\sum_{\rho\in\llbracket p\rrbracket}[\hat{{\bm{G}}}_{{\bm{T}}}]_{(i\nu),(j\rho)}V_{\mu\rho}}_{:=[\hat{{\bm{G}}}_{{\bm{T}},{\bm{V}}}]_{(i\mu),(j\nu)}}\Omega_{j\nu}$ (69)
$\implies$ $\bm{\Omega}={\rm mat}\left(\left({\bm{I}}_{np}+\hat{\bm{G}}_{{\bm{T}},{\bm{V}}}\right)^{-1}{\rm vec}\left(\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}\right)\right),$ (70)
and
$\bm{\mathbf{\hat{W}}}(\gamma{\bm{V}}-{\bm{A}}^{T})=\bm{\mathbf{X}}{\rm mat}\left(\hat{{\bm{G}}}_{\bm{\mathcal{T}}}\left({\bm{I}}_{np}+\hat{\bm{G}}_{{\bm{T}},{\bm{V}}}\right)^{-1}{\rm vec}\left(\bm{\mathbf{X}}\bm{\mathbf{\hat{W}}}\right)\right)-\bm{\mathbf{\hat{W}}}{\bm{A}}^{T}$ (71)
$\implies$ $a\gamma{\rm vec}\left(\bm{\mathbf{\hat{W}}}{\bm{V}}\right)={\bm{T}{\rm vec}\left(\bm{\mathbf{\hat{W}}}{\bm{V}}\right)},$ (72)
@@
$\mathbb{R}^{p\times p}\ni\left[\hat{{\bm{G}}}_{\bm{\mathcal{T}}}\left({\bm{I}}_{np}+\hat{\bm{G}}_{{\bm{T}},{\bm{V}}}\right)^{-1}\right]_{(i.),(j.)}=$ $\delta_{ij}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T})^{-1}$ (73)
$\left[\left(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}+{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}\right)(a{\bm{I}}-{\bm{V}}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T})^{-1}\right]^{-1}$ (74)
$=\delta_{ij}a^{-1}\bm{\mathcal{T}}({\mathrm{y}}_{i}){\bm{V}}^{T}$ (75)
Therefore, $\bm{\mathbf{\hat{W}}}=\bm{\mathbf{\hat{W}}}_{\bm{T}}{\bm{V}}^{-1}$ is a fixed point of the algorithm, for $a_{t}$ and $\gamma_{t}$ appropriately chosen, with eigenvalue given by $a_{t}\gamma_{t}$ at convergence. [^2] Note that the overlap matrices for this algorithm refer to $\bm{\mathbf{\hat{W}}}_{\bm{T}}{\bm{V}}^{-1}$ and not directly to the spectral estimator itself. However, if $\|{\bm{M}}\|_{\mathrm{F}}>0$, the weak recovery condition is satisfied.